A projectile is fired from a gun with a muzzle speed of 2000m/s at an elevation of 45degree.it maximum height would approximately be same as that reached by an identical body fired vertically with what muzzle velocity?

Grade:11

1 Answers

DIWAKAR PANDEY
6 years ago
For same height this {(2)means square} U(2)sin2x/2g=u`(2)/g 4000000/2×2×10=u`(2)/g100000=u(`2)/101000000=u`(2) U`=√1000000 U`=1000m/s